OpenStudy (anonymous):

a local toy store sells whistles in packages of 3,6,or 12. if you need 30 whistles for a party, in how many diffrent combinations can you buy them?

OpenStudy (anonymous):

hi hero

hero (hero):

Hi :D

OpenStudy (anonymous):

how r u long time no see lol

hero (hero):

I'm good. Where have you been?

OpenStudy (anonymous):

i started working and been busy like crazy i started new classes as well so been super busy

hero (hero):

I can imagine

hero (hero):

Satellite, how do you do this?

OpenStudy (anonymous):

these r so confusing its like i knkow how to do it but dont know how

OpenStudy (anonymous):

dunno i am just trying stuff out

OpenStudy (anonymous):

3*10 6*5 2*12+6 2*12+3*2 6*4+6 6*4+3*2 6*3+12 ...

OpenStudy (anonymous):

count them

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

6*3+3*4 6*2+3*6

hero (hero):

There's something about this I don't like

OpenStudy (anonymous):

like what?

hero (hero):

The "counting" part

OpenStudy (anonymous):

i wonder is there is a slick way, because i don't see it \[2\times 12+6\] \[2\times 12+ 3 + 3\] \[12+2\times 6\] \[12+6+2\times 3\] \[12+6\times 3\] \[5\times 6\] \[4\times 6+2\times 3\] \[3\times 6+4\times 3\] \[2\times 6+6\times 3\] \[1\times 6+8\times 3\] \[10\times 3\]

OpenStudy (anonymous):

still don't see a slick way but i am sure there is one

OpenStudy (anonymous):

think about it like this: there are 3 types of combinations that will yield our desired result. type 1 is where all the bags have the same amount. type 2 is where the bags have two amounts. type 3 is where the bags have 3 different amounts.

hero (hero):

iraqichick, what have you been learning in class?

OpenStudy (anonymous):

this is the first week we do these its like mind stuff

OpenStudy (anonymous):

there are 12 ways, and since that is small enough i guess you just count

OpenStudy (anonymous):

problem solving strategies

OpenStudy (anonymous):

i like the strategy of brute force for this one

OpenStudy (anonymous):

but still i dont get it its like getting more confusing with every problem!

hero (hero):

I can imagine

OpenStudy (anonymous):

so nobody knows how to solve this??

OpenStudy (anonymous):

count all the different combination that give you 30 items.

OpenStudy (anonymous):

so 11?

OpenStudy (anonymous):

i think it is 12 let me check again

OpenStudy (anonymous):

11 was wrong lol

OpenStudy (anonymous):

the way the exaple does it is they made a table but its so confusing cuz they dont tell u wat they did or how they solved. and yest its 12

OpenStudy (anonymous):

i think 12

hero (hero):

I would like to see that table

OpenStudy (anonymous):

the key to thinking about this is that the sum must be 30. In all combinatorics problems, there will be a constraint similar to this. You just have to reason through it, and count them up

hero (hero):

I would have figured it out. I just assumed there was some kind of formula associated with it

OpenStudy (anonymous):

there are some counting rules, but the sum does not allow for them to be used

OpenStudy (anonymous):

two ways if you have 2 packs of 12 they are 2 packs of 12 and 1 pack of 6 2 packs of 12 and 2 packs of 3 4 ways using on pack of 12 they are 1 pack of 12, 1 pack of 6 and 4 packs of 3 1 pack of 12, 2 packs of 6 and 2 packs of 3 1 pack of 12 and 6 packs of 3 1 pack of 12 and 3 packs of 6 then if no packs of 12 are used you can use 5 packs of six 4 packs of six 3 packs of six 2 packs of six 1 packs of six 0 packs of six total of 12

hero (hero):

Combinations is definitely not one of my favorite math topics